# SePPO: Semi-Policy Preference Optimization for Diffusion Alignment

### v1

@@ 4.1 Iterative Alignment from Semi-policy Preference Optimization
Figure 2: Ablation study of the reference model selection strategies. The x-axis is the number of diffusion steps $T$. The y-axis is the testing score. (The higher is better.)
Algorithm 1 SePPO
1:  for $k=1,\cdots,K$ do
2:          Sample a reference model ${\rm ref}\sim\mathcal{U}(0,k-1)$.
3:          Generate $\mathbf{x}_{0}^{\rm ref}$ from $\theta_{\rm ref}$, and compose data pairs $(\mathbf{x}_{0}^{w},\mathbf{x}_{0}^{\rm ref},\mathbf{c})$.
4:          $\theta_{k}\leftarrow\theta_{k-1}-\eta_{k-1}\nabla\mathcal{L}(\theta;\mathbf{x}_{0}^{w},\mathbf{x}_{0}^{\rm ref})$    # Or other optimizer, e.g., AdamW.
5:  end for
5:  ${\theta_{K}}$
4.2 Anchor-based Adaptive Flipper for Preference Optimization
On the other hand, this type of demonstrated method faces a significant problem: When using generated reference images $\mathbf{x}_{0}^{\rm ref}$, we cannot determine the relative relationship between the reference model and the current model (whether the reference image is better than the image generated by the current model or not). Existing methods, when using a reference model to generate reference samples, always directly assume that the reference samples have relatively poor quality. For example, DITTO assumes that the earlier checkpoints used as the reference model produce worse-performing samples. However, as shown in the experiment in Figure 2, the reference model is not always inferior to the current model.
Specifically, based on the DPO-style optimization method, we analyze the change in gradient updates when the reference image is better than the image generated by the current model. The gradient of the loss function in equation 6 is
$\nabla\mathcal{L}(\theta;\mathbf{x}_{0}^{w},\mathbf{x}_{0}^{\rm ref})~{}=~{}\mathop{\mathbb{E}}_{\epsilon^{w},\epsilon^{\rm ref},t}\Big{[}$ $-2\beta Tw_{t}{\sigma}\Big{(}-\beta Tw_{t}(\hat{\sigma}_{\rm ref}^{2}-\hat{\sigma}_{w}^{2})\Big{)}\underbrace{\Big{(}\epsilon_{\theta}(\mathbf{x}_{t}^{w})-\mathbf{\epsilon}^{w}-\epsilon_{\theta}(\mathbf{x}_{t}^{\rm ref})+\mathbf{\epsilon}^{\rm ref}\Big{)}}_{\text{error term}}\Big{]},$ (7)
where $\hat{\sigma}_{w}^{2}\coloneqq\|\epsilon_{\theta}(\mathbf{x}_{t}^{w})-\mathbf{\epsilon}^{w}\|^{2}-\|\epsilon_{\rm ref}(\mathbf{x}_{t}^{w})-\mathbf{\epsilon}^{w}\|^{2}$ and $\hat{\sigma}_{\rm ref}^{2}\coloneqq\|\epsilon_{\theta}(\mathbf{x}_{t}^{\rm ref})-\mathbf{\epsilon}^{\rm ref}\|^{2}-\|\epsilon_{\rm ref}(\mathbf{x}_{t}^{\rm ref})-\mathbf{\epsilon}^{\rm ref}\|^{2}$. If $\mathbf{x}_{t}^{\rm ref}$ generated by the reference model is good and close to $\mathbf{x}_{t}^{w}$, both the error term and the weight term $\sigma(\cdot)$ tend to be small. This leads to a very small gradient update without making full use of the information in the data sample, even though we know that $\mathbf{x}_{0}^{w}$ is generated by experts with high quality.
Therefore, in order to avoid the impact of the uncertainty of the reference image, we design a strategy that uses the winning image as an anchor, and evaluate the quality of the images generated by the reference model based on its performance relative to the current model on the anchor. Specifically, we design an Anchor-based Adaptive Flipper (AAF) to the loss function in equation 6 as follows:
@@ 4.2 Anchor-based Adaptive Flipper for Preference Optimization
$\mathcal{L}(\theta;\mathbf{x}_{0}^{w})~{}=~{}\mathop{\mathbb{E}}_{\epsilon^{w},\epsilon^{\rm ref},t}\Big{[}-\log{\sigma}\Big{(}$ $-\beta Tw_{t}\big{(}\|\epsilon_{\theta}(\mathbf{x}_{t}^{w})-\mathbf{\epsilon}^{w}\|^{2}-\|\epsilon_{\rm ref}(\mathbf{x}_{t}^{w})-\mathbf{\epsilon}^{w}\|^{2}$ (8)
@@
$-{\rm sign}\cdot(\|\epsilon_{\theta}(\mathbf{x}_{t}^{\rm ref})-\mathbf{\epsilon}^{\rm ref}\|^{2}-\|\epsilon_{\rm ref}(\mathbf{x}_{t}^{\rm ref})-\mathbf{\epsilon}^{\rm ref}\|^{2})\big{)}\Big{)}\Big{]},$
@@
where $\epsilon^{w},\epsilon^{\rm ref}\stackrel{{\scriptstyle\rm iid}}{{\sim}}\mathcal{N}(0,\mathbf{I})$ and $\mathbf{x}_{t}^{w,{\rm ref}}=\sqrt{\alpha^{t}}\mathbf{x}_{0}^{w,{\rm ref}}+\sqrt{1-\alpha^{t}}\epsilon^{w,{\rm ref}}$. ${\rm sign}$ is a binary variable defined as
@@
Intuitively, if the reference model has a higher probability of generating noise $\mathbf{\epsilon}^{w}$ compared to the current model, then in this situation, the reference image generated by the reference model is also more likely to be a winning image than a losing image. We formalize this claim in Theorem 4.1, which motivates the design of the criterion in equation 9.
Theorem 4.1.
@@
The proof of Theorem 4.1 is given in Appendix A.1. According to Theorem 4.1, if ${\rm sign}=-1$, it means that $\mathbf{x}_{0}^{\rm ref}$ generated by $\epsilon_{\rm ref}$ has a high probability to be better than the output of $\epsilon_{\theta}$. In this situation, $\mathbf{x}_{0}^{\rm ref}$ is of good quality and should not be rejected. The effectiveness of the AAF is further verified in the ablation study in Section 5.
@@ 5.1.1 Text-to-Image
In the text-to-image task, we test our methods based on the stable diffusion 1.5 (SD-1.5) model. Following Diffusion-DPO (Wallace et al., 2024), we use the sampled training set of the Pick-a-Pic v2 dataset (Kirstain et al., 2023) as the training dataset. Pick-a-Pic dataset is a human-annotated preference dataset for image generation. It consists of images generated by the SDXL-beta (Podell et al., 2024) and Dreamlike models. Specifically, as mentioned in Diffusion-DPO, we remove approximately $12\%$ pairs with ties and use the remaining $851,293$ pairs, which include $58,960$ unique prompts for training. We use AdamW (Loshchilov & Hutter, 2019) as the optimizer. We train our model on $8$ NVIDIA A100 GPUs with local batch size $1$, and the number of gradient accumulation steps is set to $256$. Thus, the equivalent batch size is $2048$. We train models at fixed-square resolutions. A learning rate of $5\times 10^{-9}$ is used, as we find that a smaller learning rate can help avoid overfitting. We set $\beta$ to $2000$, which stays the same in Diffusion-DPO. For evaluation datasets, we use the validation set of the Pick-a-Pic dataset, the Parti-prompt, and HPSv2, respectively. We utilize the default stable diffusion inference pipeline from Huggingface when testing. The metrics we use are PickScore, HPSv2 score, ImageReward score, and Aesthetic score.
@@ 5.1.2 Text-to-Video
To further verify that SePPO works well in text-to-video generation tasks, we test our methods based on the AnimateDiff (Guo et al., 2024). We use the training set from MagicTime (Yuan et al., 2024b) as our training set and utilize the ChronoMagic-Bench-150 (Yuan et al., 2024c) dataset as our validation set. We use LoRA (Hu et al., 2022) to train all the models at the resolution $256\times 256$ and $16$ frames are taken by dynamic frames extraction from each video. The learning rate is set to $5\times 10^{-6}$ and the training steps are set to $1000$. The metrics we use are FID (Heusel et al., 2017), LPIPS (Zhang et al., 2018), SSIM (Wang et al., 2004), PSNR (Hore & Ziou, 2010) and FVD (Unterthiner et al., 2019).
@@ 5.2.1 Analysis of Text-to-Image
To verify the effectiveness of the proposed SePPO, we compare SePPO with the SOTA methods, including DDPO (Black et al., 2024), D3PO (Yang et al., 2024), Diffusion-DPO, SPO (Liang et al., 2024b) and SPIN-Diffusion. We first report all the comparison results on the validation unique split of Pick-a-Pic dataset in Table 1. Specifically, SFTw indicates that we use the winning images from the Pick-a-Pic dataset for supervised fine-tuning. SePPOr and SePPOw indicate that we use randomly chosen images or winning images in the Pick-a-Pic dataset as the training set for SePPO. SePPO outperforms previous methods across most metrics, even those that utilize reward models during the training process, such as DDPO and SPO. Moreover, SePPO does not require the three-stage training process like SPIN-Diffusion, nor does it require the complex selection of hyperparameters. Notably, we observe that SePPO significantly improves ImageReward, which may be attributed to the fact that ImageReward reflects not only the alignment between the image and human preference but also the degree of alignment between the image and the text. In contrast, the other metrics primarily reflect the alignment between the image and human preference.
@@
Previous methods like SPIN-Diffusion used the winning images from the Pick-a-Pic dataset as training data. In our experiments, aside from training using the winning images as done previously, we also conduct an experiment where we randomly select images from both the winner and the losing sets as training data for SePPO, which we refer to as SePPOr. Despite using a lower-quality training data distribution, SePPOr still outperforms other methods on most metrics. Notably, when compared to SFTw, which is fine-tuned on the winning images, SePPOr still exceeds SFTw on three key metrics, further demonstrating the superiority of SePPO.
@@
To better evaluate SePPO’s out-of-distribution performance, we also test the model using the HPSv2 and Parti-prompt datasets, which have different distributions from the Pick-a-pic dataset. As shown in Table 2, SePPO outperforms all other models on these datasets. It is worth noting that SPO has not been tested on these two datasets and SPIN-Diffusion does not report their precise results. So, we reproduce the results by using the checkpoints of SPIN-Diffusion [^1] https://huggingface.co/UCLA-AGI/SPIN-Diffusion-iter3 and SPO [^2] https://huggingface.co/SPO-Diffusion-Models/SPO-SD-v1-5_4k-p_10ep available on HuggingFace, referred to as SPIN-Diffusion and SPO.
On the left side of Figure 3, we further explore the relationship between our AAF rate and model performance. AAF rate is defined as the ratio (#$(\#{\rm{sign}}=-1)/(\#{\rm total})$ in a batch of data. The AAF rate reflects the comparative performance between the current model and all previous checkpoints; a higher AAF rate indicates that the samples generated by the reference model are more likely to be negative samples for the current model, meaning that the current model is performing better relative to all previous models. We observe that, as training steps increase, both the AAF rate and PickScore gradually increase, showing a similar trend. We then display the changes in PickScore during the training process of SePPO and SFT on the right side of Figure 3. SFT quickly converges and begins to fluctuate, while SePPO is able to steadily improve throughout the training process.
@@
We also visualize the results of SD-1.5, SPO, SPIN-Diffusion, and SePPO in Figure 4. SePPO is able to capture the verb “nested” and also generating better eye details. SePPO not only demonstrates superior visual quality compared to other methods but also excels in image-text alignment. This is because SePPO’s training approach, which does not rely on a reward model to guide the learning direction, allows the model to learn both human preferences and improve in areas that reward models may fail to address.
Table 1: Model Feedback Results on the Pick-a-Pic Validation Set.
Methods PickScore $\uparrow$ HPSv2 $\uparrow$ ImageReward $\uparrow$ Aesthetic $\uparrow$
@@
SPIN-Diffusion 21.55 27.10 0.484 5.929
SePPOr 21.33 27.07 0.524 5.712
SePPOw 21.57 27.20 0.615 5.772
Table 2: Model Feedback Results on the HPSv2 and Parti-prompt Datasets.
PickScore $\uparrow$ HPSv2 $\uparrow$ ImageReward $\uparrow$ Aesthetic $\uparrow$
Methods HPS Parti HPS Parti HPS Parti HPS Parti
SD-1.5 20.95 21.38 27.17 26.70 0.08 0.16 5.55 5.33
SFTw 21.50 21.68 27.88 27.40 0.68 0.56 5.82 5.53
Diffusion-DPO 21.40 21.63 27.23 26.93 0.30 0.32 5.68 5.41
Diffusion-RPO 21.43 21.66 27.37 27.05 0.34 0.39 5.69 5.43
SPO 21.87 21.85 27.60 27.41 0.41 0.42 5.87 5.63
SPIN-Diffusion 21.88 21.91 27.71 27.58 0.54 0.51 6.05 5.78
SePPOw 21.90 21.93 27.92 27.69 0.70 0.58 5.94 5.64
(a) AAF Rate and PickScore
(b) SFT vs. SePPO
Figure 3: Relations of AAF rates and model performance.
Figure 4: Text-to-image generation results of SD-1.5, SPO, SPIN-Diffusion and SePPO. The prompts from left to right are: (1) Photo of a pigeon in a well tailored suit getting a cup of coffee in a cafe in the morning; (2) Ginger Tabby cat watercolor with flowers; (3) An image of a peaceful mountain landscape at sunset, with a small cabin nestled in the trees and a winding river in the foreground; (4) Space dog; (5) b&w photo of 42 y.o man in white clothes, bald, face, half body, body, high detailed skin, skin pores, coastline, overcast weather.
Figure 5: Text-to-video generation results of AnimateDiff (Raw), SFT, and SePPO. The prompt is: “Time-lapse of a lettuce seed germinating and growing into a mature plant. Initially, a seedling emerges from the soil, followed by leaves appearing and growing larger. The plant continues to develop…

### v2

@@ 4.1 Random Checkpoint Replay
Figure 2: Ablation study of the Experience Replay Distribution selection strategies. The x-axis is the number of training steps $T$. The y-axis is the testing score. (The higher is better.)
We then use ERD to generate the pseudo “losing samples” required for the DPO training process. Thus, in each iteration, after omitting $\mathbf{c}$ for the sake of concision, the loss function adapted from diffusion DPO is as follows
@@
$\mathcal{L}$ $(\theta;\mathbf{x}_{0}^{w},\mathbf{x}_{0}^{\rm rand})~{}=~{}\mathop{\mathbb{E}}_{\epsilon^{w},\epsilon^{\rm ref},t}\Big{[}-\log\cdot$ (5)
${\sigma}$ $\Big{(}-\beta Tw_{t}(\|\epsilon_{\theta}(\mathbf{x}_{t}^{w})-\mathbf{\epsilon}^{w}\|^{2}-\|\epsilon_{\rm ref}(\mathbf{x}_{t}^{w})-\mathbf{\epsilon}^{w}\|^{2}-\|\epsilon_{\theta}(\mathbf{x}_{t}^{\rm rand})-\mathbf{\epsilon}^{\rm ref}\|^{2}+\|\epsilon_{\rm ref}(\mathbf{x}_{t}^{\rm rand})-\mathbf{\epsilon}^{\rm ref}\|^{2})\Big{)}\Big{]},$
where $t\sim\mathcal{U}(1,T)$, $\epsilon^{w},\epsilon^{\rm ref}\stackrel{{\scriptstyle\rm iid}}{{\sim}}\mathcal{N}(0,\mathbf{I})$, and $\mathbf{x}_{t}^{w,{\rm rand}}=\sqrt{\alpha^{t}}\mathbf{x}_{0}^{w,{\rm rand}}+\sqrt{1-\alpha^{t}}\epsilon^{w,{\rm ref}}$. Notably, in equation 5, the reference image $\mathbf{x}_{0}^{\rm rand}$ is generated from the ERD distribution, which is on-policy learning in the last two settings.
We use the Pick-a-Pic dataset and the stable diffusion 1.5 (SD-1.5) model to conduct all experiments for this study. In experiments, we save a checkpoint every $30$ updates ($K=7$ iterations). The experimental results are shown in Figure 2. In the beginning, the performances of all models are improved. However, as the number of steps $T$ increases, the performance of the model with ${\rm ERD}=0$ starts to decrease and quickly exhibits overfitting. The model with ${\rm ERD}=k-1$ shows unstable performances. Due to the fact that when all “losing samples” are generated from the last checkpoint in the current round of training, the model is prone to fall into non-global optima, leading to instability in the training process Florensa et al. (2018). Therefore, considering both model performance and stability, our method uses ${\rm ERD}=[0,k-1]$ to generate reference samples, referring to this approach as Random Checkpoint Replay.
Specifically, in Theorem 4.1, under the PAC-Bayesian framework Germain et al. (2016); Seeger (2002), we theoretically show that our uniformly sampling strategy has better generalization ability compared to the latest checkpoint selection strategy, e.g., SPIN.
Theorem 4.1
With uniformly sampling from previous checkpoints $[0,K-1]$, the PAC-Bayesian upper bound of the generalization loss is guaranteed to be smaller or equal to the latest checkpoint.
$\mathcal{O}\Big{(}\mathop{\mathbb{E}}_{\pi_{\theta}\sim Q_{\rm UNI}}\big{[}{L}(\pi_{\theta})\big{]}\Big{)}\leq\mathcal{O}\Big{(}\mathop{\mathbb{E}}_{\pi_{\theta}\sim Q_{K}}\big{[}{L}(\pi_{\theta})\big{]}\Big{)},$ (6)
where ${L}(\pi_{\theta})\coloneqq\mathop{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}}\mathcal{L}(\theta;\mathbf{x})$ is the generalization loss on the policy, $Q_{\rm UNI}$ is the posterior policy distribution with uniformly sampling, and $Q_{K}$ is the posterior distribution with the latest checkpoint.
The complete proof is given in Appendix A.1.
4.2 Self-Sampling Regularization
On the other hand, is the data sampled from ERD truly “losing samples” relative to the current policy? Obviously, when generating the samples $\mathbf{x}_{0}^{\rm rand}$, we cannot determine the relative relationship between the ERD and the current model (whether the image sampled from ERD is better than the image generated by the current model or not). Existing methods always directly assume that the ERD samples have relatively poor quality. For example, DITTO assumes that the earlier checkpoints used as the reference model produce worse-performing samples. However, as shown in the experiment in Figure 2, the ERD is not always inferior to the current model.
Specifically, based on the DPO-style optimization method, we analyze the change in gradient updates when the ERD image is better than the image generated by the current model. The gradient of the loss function in equation 5 is as follows
$\nabla\mathcal{L}(\theta;\mathbf{x}_{0}^{w},\mathbf{x}_{0}^{\rm rand})~{}=~{}\mathop{\mathbb{E}}_{\epsilon^{w},\epsilon^{\rm ref},t}\Big{[}-2\beta Tw_{t}{\sigma}\Big{(}-\beta Tw_{t}(\hat{\sigma}_{\rm ref}^{2}-\hat{\sigma}_{w}^{2})\Big{)}\underbrace{\Big{(}\epsilon_{\theta}(\mathbf{x}_{t}^{w})-\mathbf{\epsilon}^{w}-\epsilon_{\theta}(\mathbf{x}_{t}^{\rm rand})+\mathbf{\epsilon}^{\rm ref}\Big{)}}_{\text{error term}}\Big{]},$ (7)
where $\hat{\sigma}_{w}^{2}\coloneqq\|\epsilon_{\theta}(\mathbf{x}_{t}^{w})-\mathbf{\epsilon}^{w}\|^{2}-\|\epsilon_{\rm ref}(\mathbf{x}_{t}^{w})-\mathbf{\epsilon}^{w}\|^{2}$ and $\hat{\sigma}_{\rm ref}^{2}\coloneqq\|\epsilon_{\theta}(\mathbf{x}_{t}^{\rm rand})-\mathbf{\epsilon}^{\rm ref}\|^{2}-\|\epsilon_{\rm ref}(\mathbf{x}_{t}^{\rm rand})-\mathbf{\epsilon}^{\rm ref}\|^{2}$. If $\mathbf{x}_{t}^{\rm rand}$ generated by the ERD is good and close to $\mathbf{x}_{t}^{w}$, both the error term and the weight term $\sigma(\cdot)$ tend to be small. This leads to a very small gradient update without making full use of the information in the data sample, even though we know that $\mathbf{x}_{0}^{w}$ is generated by experts with high quality.
Therefore, in order to avoid the impact of the uncertainty of the image sampled from ERD, we design a strategy that uses the winning image as an anchor, and evaluates the quality of the ERD samples based on its performance relative to the current model on the anchor. Specifically, we design the Self-sampling Regularization (SSR) to the loss function in equation 5 as follows:
@@ 4.2 Self-Sampling Regularization
$\mathcal{L}(\theta;\mathbf{x}_{0}^{w},\mathbf{x}_{0}^{\rm rand})~{}=~{}\mathop{\mathbb{E}}_{\epsilon^{w},\epsilon^{\rm ref},t}\Big{[}-\log\cdot$ (8)
${\sigma}\Big{(}-\beta Tw_{t}\big{(}\|\epsilon_{\theta}(\mathbf{x}_{t}^{w})-\mathbf{\epsilon}^{w}\|^{2}-\|\epsilon_{\rm ref}(\mathbf{x}_{t}^{w})-\mathbf{\epsilon}^{w}\|^{2}-{\rm sign}\cdot(\|\epsilon_{\theta}(\mathbf{x}_{t}^{\rm rand})-\mathbf{\epsilon}^{\rm ref}\|^{2}-\|\epsilon_{\rm ref}(\mathbf{x}_{t}^{\rm rand})-\mathbf{\epsilon}^{\rm ref}\|^{2})\big{)}\Big{)}\Big{]},$
@@
where $\epsilon^{w},\epsilon^{\rm ref}\stackrel{{\scriptstyle\rm iid}}{{\sim}}\mathcal{N}(0,\mathbf{I})$ and $\mathbf{x}_{t}^{w,{\rm rand}}=\sqrt{\alpha^{t}}\mathbf{x}_{0}^{w,{\rm rand}}+\sqrt{1-\alpha^{t}}\epsilon^{w,{\rm ref}}$. ${\rm sign}$ is a binary variable defined as
@@
Bridging SFT and DPO. We give the pseudocode of SPPO in Algorithm 1. The integration of our random checkpoint replay and self-sampling regularization strategies facilitates seamless switching between SFT and DPO, by leveraging results from randomly sampled checkpoints. Specifically in Line 4 of Algorithm 1, the loss function is calculated using equation 8, where function ${\rm sign}$ judges whether SFT or DPO is performed. If ${\rm sign}=1$, it performs DPO updates with $\mathbf{x}_{0}^{w}$ and $\mathbf{x}_{0}^{{\rm rand}_{w}}$ targets. If ${\rm sign}=-1$, the performance is equivalent to SFT updates with $\mathbf{x}_{0}^{w}$ and $\mathbf{x}_{0}^{{\rm rand}_{l}}$ labels, because the $\log\sigma(\cdot)$ function is monotonic. We validate the effectiveness of SSR through the ablation study in Table 6 in Section 5.2.4 and show the SSR Rate in Figure 4 where the definition of SSR Rate is given in Section 5.2.1.
Intuitively, if the sampled checkpoint has a higher probability of generating noise $\mathbf{\epsilon}^{w}$ compared to the current model, then in this situation, the reference image generated by the sampled checkpoint is also more likely to be a winning image than a losing image. We formalize this claim in Theorem 4.2, which motivates the design of the criterion in equation 9.
Theorem 4.2
@@
The proof of Theorem 4.2 is given in Appendix A.2. According to Theorem 4.2, if ${\rm sign}=-1$, it means that $\mathbf{x}_{0}^{\rm rand}$ generated by $\epsilon_{\rm ref}$ has a high probability to be better than the output of $\epsilon_{\theta}$. In this situation, $\mathbf{x}_{0}^{\rm rand}$ is of good quality and should not be rejected. The effectiveness of the SSR is further verified in the ablation study in Section 5.
Algorithm 1 SSPO
1:  for $k=1,\cdots,K$ do
2:          Sample a checkpoint ${\rm rand}\sim\mathcal{U}(0,k-1)$.
3:          Generate $\mathbf{x}_{0}^{\rm rand}$ from $\theta_{\rm rand}$, and compose data pairs $(\mathbf{x}_{0}^{w},\mathbf{x}_{0}^{\rm rand},\mathbf{c})$.
4:          Compute $\mathcal{L}(\theta;\mathbf{x}_{0}^{w},\mathbf{x}_{0}^{\rm rand})$ according to equation 8.
5:          $\theta_{k}\leftarrow\theta_{k-1}-\eta_{k-1}\nabla\mathcal{L}(\theta;\mathbf{x}_{0}^{w},\mathbf{x}_{0}^{\rm rand})$    # Or other optimizer, e.g., AdamW.
6:  end for
6:  ${\theta_{K}}$
@@ 5.1.1 Text-to-Image
In the text-to-image task, we test our methods based on the stable diffusion 1.5 (SD-1.5) model. Following Diffusion-DPO (Wallace et al., 2024), we use the sampled training set of the Pick-a-Pic v2 dataset (Kirstain et al., 2023) as the training dataset. Pick-a-Pic dataset is a human-annotated preference dataset for image generation. It consists of images generated by the SDXL-beta (Podell et al., 2024) and Dreamlike models. Specifically, as mentioned in Diffusion-DPO, we remove approximately $12\%$ pairs with ties and use the remaining $851,293$ pairs, which include $58,960$ unique prompts for training. We use AdamW (Loshchilov & Hutter, 2019) as the optimizer. We train our model on $8$ NVIDIA A100 GPUs with local batch size $1$, and the number of gradient accumulation steps is set to $256$. Thus, the equivalent batch size is $2048$. A learning rate of $5\times 10^{-9}$ is used, as we find that a smaller learning rate can help avoid overfitting. We set $\beta$ to $2000$, which stays the same in Diffusion-DPO. For evaluation datasets, we use the validation set of the Pick-a-Pic dataset, the Parti-prompt, and HPSv2, respectively. We utilize the default stable diffusion inference pipeline from Huggingface when testing. The metrics we use are PickScore, HPSv2, ImageReward, and Aesthetic score.
@@ 5.1.2 Text-to-Video
To further verify that SSPO works well in text-to-video generation tasks, we test our methods based on the AnimateDiff (Guo et al., 2024). We use the training set from MagicTime (Yuan et al., 2024b) as our training set and utilize the ChronoMagic-Bench-150 (Yuan et al., 2024c) dataset as our validation set. We use LoRA (Hu et al., 2022) to train all the models at the resolution $256\times 256$ and $16$ frames are taken by dynamic frames extraction from each video. The learning rate is set to $5\times 10^{-6}$ and the training steps are set to $1000$. The metrics we use are FID (Heusel et al., 2017), PSNR (Hore & Ziou, 2010) and FVD (Unterthiner et al., 2019), MTScore (Yuan et al., 2024c) and CHScore(Yuan et al., 2024c).
@@ 5.2.1 Analysis of Text-to-Image
To verify the effectiveness of the proposed SSPO, we compare SSPO with the SOTA methods, including DDPO (Black et al., 2024), DPOK (Fan et al., 2023), D3PO (Yang et al., 2024), Diffusion-DPO, Diffusion-RPO (Gu et al., 2024), SPO (Liang et al., 2024b) and SPIN-Diffusion. We first report all the comparison results on the validation unique split of Pick-a-Pic dataset in Table 1. SSPO outperforms previous methods across most metrics, even those that utilize reward models during the training process, such as DDPO and SPO. Moreover, SSPO does not require the three-stage training process like SPIN-Diffusion, nor does it require the complex selection of hyperparameters. Notably, we observe that SSPO significantly improves ImageReward, which may be attributed to the fact that ImageReward reflects not only the alignment between the image and human preference but also the degree of alignment between the image and the text. In contrast, the other metrics primarily reflect the alignment between the image and human preference.
@@
Previous methods like SPIN-Diffusion used the winning images from the Pick-a-Pic dataset as training data. In our experiments, aside from training using the winning images as done previously, we also conduct an experiment where we randomly select images from both the winner and the losing sets as training data for SSPO, which we refer to as SSPOr. Despite using a lower-quality training data distribution, SSPOr still outperforms other methods on most metrics. Notably, when compared to SFT, which is fine-tuned on the winning images, SSPOr still exceeds SFT on three key metrics, further demonstrating the superiority of SSPO.
@@
To better evaluate SSPO’s out-of-distribution performance, we also test the model using the HPSv2 and Parti-prompt datasets, which have different distributions from the Pick-a-pic dataset. As shown in Table 2 and Table 7 in the Appendix, SSPO outperforms all other models on these datasets. It is worth noting that SPO has not been tested on these two datasets and SPIN-Diffusion does not report their precise results. So, we reproduce the results by using the checkpoints of SPIN-Diffusion [^1] https://huggingface.co/UCLA-AGI/SPIN-Diffusion-iter3 and SPO [^2] https://huggingface.co/SPO-Diffusion-Models/SPO-SD-v1-5_4k-p_10ep available on HuggingFace.
We also evaluate SSPO on the Pick-a-Pic validation set using the SDXL backbone to assess its performance in a more advanced generation setting. As shown in Table 3, SSPO consistently outperforms all baseline methods on PickScore, HPSv2, and Image Reward, indicating stronger alignment with human preference and learned reward signals. Although SPO achieves a slightly higher Aesthetic score, SSPO demonstrates superior overall performance across key metrics. These results further validate the effectiveness of SSPO in enhancing model feedback quality, even when applied to stronger base models like SDXL.
In Figure 4, we further explore the relationship between our SSR rate and model performance. SSR rate is defined as the ratio (#$(\#{\rm{sign}}=-1)/(\#{\rm total})$ in a batch of data. The SSR rate reflects the comparative performance between the current model and all previous checkpoints; A higher SSR rate indicates that the samples generated by the ERD are more likely to be positive samples compared to the current model, meaning that the current model is competitive with ERD which relative to all previous models. We observe that in the early stages of training, a low SSR rate enables the model to quickly learn information from positive samples. In the later stages, as the SSR rate increases, the current model no longer holds an absolute advantage, indicating that the model has effectively learned the distribution of the sample data.
@@
We also visualize the results of SD-1.5, SPO, SPIN-Diffusion, and SSPO in Figure 5. SSPO is able to capture the verb “nested” and also generates better eye details. SSPO not only demonstrates superior visual quality compared to other methods but also excels in image-text alignment. This is because SSPO’s training approach, which does not rely on RMs to guide the learning direction, allows the model to learn both human preferences and improve in areas that RMs may fail to address. We further conduct human evaluation experiments, with the experimental details and results provided in the Appendix B.3.
Table 1: Model Feedback Results on the Pick-a-Pic Validation Set. SSPOr indicate that we use randomly chosen images in the Pick-a-Pic dataset as the training set. IR indicates Image Reward.
Methods PickScore $\uparrow$ HPSv2 $\uparrow$ IR $\uparrow$ Aesthetic $\uparrow$
@@
SPIN-Diff 21.55 27.10 0.484 5.929
SSPOr 21.33 27.07 0.524 5.712
SSPO 21.57 27.20 0.615 5.772
Table 2: Model Feedback Results on the HPSv2 Dataset. IR indicates Image Reward.
Methods PickScore $\uparrow$ HPSv2 $\uparrow$ IR $\uparrow$ Aesthetic $\uparrow$
SD-1.5 20.95 27.17 0.08 5.55
SFT 20.95 27.88 0.68 5.82
Diff-DPO 20.95 27.23 0.30 5.68
Diff-RPO 21.43 27.37 0.34 5.69
SPO 21.87 27.60 0.41 5.87
SPIN-Diff 21.88 27.71 0.54 6.05
SSPO 21.90 27.92 0.70 5.94
Table 3: Model Comparisons on SDXL on the Pick-a-Pic Validation Set. IR indicates Image Reward.
Methods PickScore $\uparrow$ HPSv2 $\uparrow$ IR $\uparrow$ Aesthetic $\uparrow$
SDXL 21.95 26.95 0.54 5.95
Diff-DPO 22.64 29.31 0.94 6.02
MAPO 22.11 28.22 0.72 6.10
SPO 23.06 31.80 1.08 6.36
SSPO 23.24 32.04 1.16 6.32
Figure 3: Correlation between SSR Rate and PickScore. There is a significant correlation between SSR Rate and PickScore.
Figure 4: Ablation study on checkpoint saving frequency. The checkpoint is saved every k steps. ($K=10/20/50$)
Figure 5: Text-to-image generation results of SD-1.5, SPO, SPIN-Diffusion and SSPO. The prompts from left to right are: (1) Photo of a pigeon in a well tailored suit getting a cup of coffee in a cafe in the morning; (2) Ginger Tabby cat watercolor with flowers; (3) An image of a peaceful mountain landscape at sunset, with a small cabin nestled in the trees and a winding river in the foreground; (4) Detailed Portrait Of A Disheveled Hippie Girl With Bright Gray Eyes By Anna Dittmann, Digital Painting, 120k, Ultra Hd, Hyper Detailed, Complimentary Colors, Wlop, Digital Painting; (5) b&w photo of 42 y.o man in white clothes, bald, face, half body, body, high detailed skin, skin pores, coastline, overcast weather.
5.2.2 Memory and Time Cost
We show the comparison of GPU and Time cost as well as performance comparisons in Table  4. We set the batch size to $1$ and calculated the average time cost and maximum GPU usage over $100$ data points for each method. SSPO is comparable to SPIN, DPOK, etc., in terms of GPU usage and execution time, second only to DPO. This is because our method does not require any reward model, eliminating the need for GPU usage and inference time of a reward model. Compared to other methods, our approach also requires fewer hyperparameters. In SSPO, compared to Diffusion-DPO, the only hyperparameter introduced is the number of iterations $K$ where we study the ablation in Figure. 4.
Table 4: Memory & time cost, Extra Hyper-parameters compared to Diff-DPO, and performance comparisons. Ex Hype. indicates Extra Hyper-parameters compared to Diff-DPO.
Methods GPU (MB) Time (s) Ex Hype.
Diff-DPO 11,206 1.36 0
D3PO 17,850 18.17 3
DPOK 16,248 9.80 2
SPIN-Diff 17,284 2.65 4
SSPO 17,284 2.65 1